# Finding Phones Fast: Low-Latency and Scalable Monitoring of Cellular Communications in Sensitive Areas

### v2

@@ Abstract.
We propose WaveTag, the first low-latency, operator-independent and scalable system designed to monitor 5G and LTE connections across all operators prior to any user data transmission. WaveTag consists of several downlink sniffers and a distributed network of uplink sniffers that measure both downlink protocol information and uplink signal characteristics at multiple locations to gain a detailed spatial image of uplink signals. WaveTag then aggregates the recorded information, processes it, and provides a decision about the connection–all done prior to the complete connection establishment of a UE. To evaluate WaveTag, we deployed it in the context of geofencing, where WaveTag was able to determine whether the signals originate from inside or outside of an area within 2.3 ms of the initial base station-to-device message, therefore enabling prompt and targeted suppression of communication before any user data was transmitted. WaveTag achieved 99.66% geofencing classification accuracy, using only the information collected before connection establishment. Finally, we conduct a real-world uplink measurement evaluation on a commercial 5G SA network.
@@ 1. Introduction
Cellular operators could provide such monitoring solutions (cherian_lte_2013, ; margolies_can_2017, ). However, several reasons make this impractical: (1) Accuracy and Cost: Under the current LTE specification (noauthor_lte_nodate, ; noauthor_stage_nodate, ), operators cannot monitor a user’s connection with the precision required in these situations (cherian_lte_2013, ; mahyuddin_overview_2017, ), especially in NLoS scenarios involving many indoor locations. Improving precision around sensitive areas would require deploying a much denser base-station infrastructure than exists today. (2) Latency: No 3GPP-specified solution provides a geofencing result based solely on initial uplink messages; localization happens later in the connection. Without a specification change, an operator aiming to prevent communication from sensitive areas would need to block all communication until geofencing results arrive. This would significantly degrade user experience for everyone near sensitive areas, who would have to wait before making calls or sending messages—an unacceptable impact in urban environments. (3) Feasibility: Every operator in the country would need to deploy this solution and collaborate with sensitive-area administrators. If even one operator refuses or is unable to cooperate, an adversary could bypass the system through that operator. Such nationwide coordination seems infeasible in practice.
On the other hand, state-of-the-art research (kotuliak_ltrack_2022, ; oh_enabling_2024, ; kumar_lte_2014, ) explores operator-independent monitoring but faces practical limitations. (1) None shows to be low-latency when deployed in a distributed system. (2) They are restricted to a single cell per SDR, whereas realistic deployments involve multiple cells across different operators. (3) They combine downlink and uplink receivers into a single device, limiting their use to locations with good downlink visibility. These challenges make it difficult to collect high-quality signal metrics and integrate them with geofencing systems. (4) They do not support 5G SA.
@@
  • We build a new data collection platform for collecting connection data based on COTS modems and collect a huge dataset of >170,000 connections with all the signal features for the initial UL messages. We make this dataset public.
@@ 2.2. Physical Uplink Channels
On the physical layer, a UE sends a message in a resource grid consisting of OFDM symbols with multiple subcarriers in each symbol. Subcarrier spacing if 15kHz for all the monitored FDD bands, and 30kHz for the 5G TDD band. For 15kHz subcarrier spacing, 14 OFDM symbols form a 1 millisecond long subframe, and 10 subframes form one frame. For 30kHz subcarrier spacing, one subframe consists of 28 OFDM symbols instead. One physical resource block (PRB) spans 12 subcarriers over 7 OFDM symbols.
@@ 3. System Design
Figure 3. In (A), we show a sketch of a deployment with two downlink receivers and six uplink receivers positioned around the sensitive area. The purple triangles on the uplink receivers show the azimuth of the two opposite facing antennas. In (B), we show how the system components interact. First, the downlink receivers detect a new connection on the downlink and inform the UL receivers about it. The UL receivers then measure the corresponding uplink signals. Finally, all measurements are aggregated in the central unit (CU), which makes a decision about the connection. In (C), we show how a concrete wall can be used for improved inside/outside separation of a single uplink receiver unit.
@@
Accurate spatial inference about a UE connection, such as precise localization or geofencing, typically requires multiple observations of uplink messages. Given the limited number of uplink messages available before connection establishment with the core network, WaveTag overcomes this constraint by deploying a distributed set of uplink receivers strategically positioned in and around the sensitive area.
@@ 3.1. System Model and Assumptions
Due to the requirement for UE detection and decision before full connection establishment with the core network, WaveTag collects metrics only for messages sent by the UE before full connection establishment. This includes at most one PRACH preamble, one RRC Connection Request on PUSCH, and one or two subsequent PUCCH messages. Subsequent PUSCH messages after the RRC Connection Request could already include messages such as a Service Request, reestablishing an existing security context and effectively communicating with the core network.
@@ 3.3. Uplink Receivers
Uplink receivers, operating also on an SDR-based platform, are connected to two wide-beam directional antennas facing away from each other, as pictured in Figure 3. This antenna orientation enables uplink receivers to estimate the direction of the signal (see Section 5 for more details). Each SDR streams IQ samples of the uplink of one or more cellular frequency bands to an uplink receiver, where it is buffered for at least 20 milliseconds. One of the main advantages of WaveTag over existing solutions is that uplink receivers are decoupled from downlink receivers. The uplink receivers do not need to monitor downlink frequencies, and therefore they do not need to acquire a sync to the xNBs or decode downlink messages. Instead, they detect and measure uplink signals using the correct uplink allocations and timings received from the downlink receiver.
@@
All relevant downlink control messages, including uplink allocations from RARs, are sent to all uplink receivers over a low-latency communication link. Each uplink receiver then independently looks up the corresponding recorded time and frequency resources for the uplink message in the buffered IQ stream. WaveTag then extracts the signal characteristics (see Section 5) based on the reference signals and finally sends them to the central unit.
Using this process, uplink receivers detect and measure all three types of uplink messages in LTE: PRACH, PUSCH, and PUCCH. In 5G, only PRACH and PUSCH feature extraction is currently implemented.
@@ 3.4. Central Unit
Measurements of individual uplink messages are sent over a low-latency communication link and aggregated at the central unit. Once we receive measurements from every uplink receiver, we can further process the signal metrics for geofencing (or localization) purposes.
@@ 4. WaveTag Implementation
In this section, we explain how our system achieves both low latency and real-world scalability, including monitoring multiple cells per operator across all their allocated frequency bands. To this end, we show how the radio units acquire and process IQ samples and how the sub-components communicate with each other.
@@ 4.1. Downlink Monitoring
We note that while downlink decoding plays a role in this process, it is not the main contribution of this paper and will not be discussed in detail. There are many downlink receiver implementations already available (erni_adaptover_2022, ; erni_glados_2025, ; falkenberg_falcon_2019, ; bui_owl_2016, ; ross_towards_2023, ), and we use XYZ [^1] To avoid de-anonymization, we will include a citation to our downlink receiver in the camera ready paper if the paper is accepted. as a base for our downlink sniffer.
The main feature of this downlink receiver is the ability to decode multiple cells in parallel on a single SDR device. It continuously tracks multiple cells from the same operator or different operators in parallel. This is achieved by splitting the single IQ stream from the SDR into multiple parallel streams using our software channelizer implementation. Our implementation is an efficient implementation of overlap save filtering  (borgerding_turning_2006, ), using FFTW3 and AVX instructions.
@@ 4.2. Uplink Processing
Classically (kotuliak_ltrack_2022, ; hoang_ltesniffer_2023, ), uplink monitoring of an LTE cell is achieved by tuning an SDR receiver to the center frequency of the uplink of the cell and collecting 30.72M IQ samples per second. However, in each of the frequency bands, there are three to four operators, each transmitting on different frequencies. Using the classical approach, we would require up to 24 SDRs in each uplink receiver to monitor the entire uplink spectrum. Such a system would be extremely expensive and difficult to maintain.
In our solution, we use SDRs with a sampling rate of at least 122.88Msps, such that they have enough bandwidth to receive a whole FDD frequency band. For the 5G TDD band with cells having as much as 100MHz bandwidth, multiple SDRs are needed. To ensure that the DC component does not disturb the uplink frequencies, we set the center frequency of our SDR between two operators allocated within a band, as close to the middle of the frequency band as possible. An uplink receiver continuously receives IQ samples from all SDRs and writes them into circular buffers. We also keep the timestamp of the last written sample. Using this timestamp, we can compute the absolute time of IQ samples in the buffer.
@@ 4.3. Message Exchange
The downlink receiver, the uplink receivers, and the central unit continuously communicate with each other over a low-latency link. They are connected with fiber links in a star architecture with a switch in the middle. Given the relatively small number of devices, the overhead of such a network topology is small. In Section 6.1, we evaluate how long a message exchange takes.
@@ 5. Low-latency Geofencing
The main challenge of precise geofencing is classifying a connection’s origin–i.e., determining whether it is inside or outside a sensitive area–based on as few uplink messages as possible. To achieve this, WaveTag uses a combination of relative signal strength and signal quality features. These features are collected from multiple uplink receivers, each equipped with two directional antennas facing in opposite directions. Finally, a classification model running on the central unit infers whether the connection originates inside or outside the perimeter.
@@ 5.2. Geofencing Model
Our classification model is a fully connected neural network that uses relative features as input. We first train the model offline with Python and pyTorch (noauthor_pytorch_nodate, ). We use a simple architecture to prevent overfitting and ensure fast inference time. The architecture consists of BatchNorm, Dropout and two fully connected layers followed by an ReLU activation function. Finally, we pass the output neuron’s result through a sigmoid function to produce the classification decision. We use binary cross-entropy as a loss function and optimize the model using an Adam optimizer. We then train the model in batches and stop training once the loss does not improve across several epochs. Finally, we perform a hyperparameter search for each deployment separately to ensure an optimal solution.
@@ 5.3. Data Collection
To improve the efficiency of the data collection, we fitted a backpack with multiple such modems that connect to the network almost every second. We can then walk a route and easily gather large amounts of uplink signal features from any location. Such an annotated dataset is crucial for training and testing of the model. We make this data available online to support further research.
@@ 6.1. Latency
In this section, we evaluate the latency of individual processing steps and also the end-to-end (E2E) latency of our system for various types of messages. Both individual steps and E2E latencies are shown in Figure 4. We use the Perfetto (google_googleperfetto_2024, ) and Tracy (taudul_wolfpldtracy_2025, ) libraries to measure the processing times of individual components during the runtime of the whole system. The exception is the measurement for the ZMQ message round-trip times, for which we built our own measurement tool.
@@
[…]
@@
In Table 1, we show the measurement results of the individual parts of the system, as well as the E2E latencies for various message types. For the E2E PRACH measurement, we measure the elapsed time from the reception of the RAR by the downlink receiver to the completion of the inference of the geofencing model. This time encapsulates the downlink decoding of the RAR, the message exchange between the downlink receiver and the uplink receiver, the PRACH measurement by the uplink receiver, the message exchange between the uplink receiver and the central unit, and finally the model inference. For the E2E PUSCH and E2E PUCCH measurements, we measure the time from the reception of the uplink message to the inference result. This measurement does not encapsulate downlink decoding and initial message exchange, as WaveTag knows about the uplink message already 4 subframes prior to the uplink transmission.
@@ 6.3.2. Second Deployment
We trained a geofencing model as a fully connected neural network model, and furthermore, we trained a logistic regression that aggregates results over multiple messages. The estimated weights from the logistic regression are: 0.39 for PRACH, 0.38 for PUSCH and 0.23 for PUCCH, approximately matching the weights from the first deployment. The accuracy of the neural network model is 99.02% on individual messages. The ensemble model then increases the accuracy for the entire connection to 99.66%. In Table 3, we report more detailed results for the final geofencing model. Throughout the deployment, we also conducted regular spot-checks with a large number of commercial phones. We did not observe any degradation in performance for any particular model.
Figure 7. Effect of uplink receivers being switched off.
Furthermore, we performed robustness tests to determine how well the model behaves if fewer uplink receivers are available or if fewer messages are available for classification. For the first test, we plot these results in Figure 7. In the second test, with just one message (PRACH), the model’s accuracy is 99.17%, and with the first two messages (PRACH and PUSCH (RRC Connection Request)), the model’s accuracy improves to 99.6%.
6.4. 5G-SA Feasibility Study
@@
With this experiment, we confirm that feature extraction works in 5G SA. Moreover, at the farthest points from the uplink receiver (1̃00m away), we still measured an SNR of more than 20dB. Furthermore, we can see a clear separation between the points on the left and right side of the field in Figure 8, meaning that the ratio of signal strengths measured on the individual ports correlates with the uplink signal direction. Given the similarities in the PHY layer of LTE and 5G, and the successful tests in the field, we conclude that if 5G SA were deployed at one of our sites, WaveTag would also correctly geofence 5G SA connections.
@@ 7. Related Work
Other projects that build operator-independent monitoring systems and rely on uplink signal measurements include (kotuliak_ltrack_2022, ; oh_enabling_2024, ; kumar_lte_2014, ). They all choose to couple uplink and downlink processing on a single host using one (or multiple) SDRs. This approach comes with two disadvantages: (1) Colocation of downlink and uplink sniffers: This constraint requires downlink sniffers to be placed only where good downlink reception is available. As a result, uplink receivers cannot be positioned indoors or in complex non-line-of-sight environments. (2) One-to-one requirement between uplink and downlink receivers: Each uplink receiver must have a dedicated downlink receiver, increasing the number of required devices and the system’s processing load. In contrast, our system decouples the downlink and uplink receivers, allowing each to be placed optimally. This design also improves scalability, as a single downlink receiver can serve multiple uplink receivers.
@@
However, the context in which these systems were implemented is also different. Mainly, they were designed to accurately localize users from multiple observations during a connection, e.g., looking for an uncooperative cellular device from a moving police car. Because to their use cases, it is enough to only observe a large number of messages on PUSCH, use rotational antennas to slowly but accurately infer the direct signal path, or potentially use active power-boosting attacks to increase the UE’s power (oh_enabling_2024, ). As a result, localization in (oh_enabling_2024, ) takes around 2 minutes, (kotuliak_ltrack_2022, ) measures ToA values of all messages from the whole connection, and (kumar_lte_2014, ) requires offline processing of recoded samples. Even though they achieve their respective tasks, these points show that they are unsuitable for low-latency geofencing. WaveTag compensates for the low number of observations with a higher number of uplink receivers, while also measuring PUCCH and PRACH messages.

### v3

@@ Abstract.
We propose WaveTag, the first low-latency, operator-independent, and scalable system designed to monitor 5G and LTE connections across all operators prior to any user data transmission. WaveTag consists of several downlink receivers and a distributed network of uplink receivers that measure both downlink protocol information and uplink signal characteristics at multiple locations to gain a detailed spatial image of uplink signals. WaveTag then aggregates the recorded information, processes it, and provides a decision about the connection before the UE completes connection establishment. To evaluate WaveTag, we deployed it in the context of geofencing, where WaveTag was able to determine whether the signals originate from inside or outside of an area within 2.3 ms of the initial base station-to-device message, therefore enabling prompt and targeted suppression of communication before any user data was transmitted. WaveTag achieved 99.66% geofencing classification accuracy, using only the information collected before connection establishment. Finally, we conduct a real-world uplink measurement evaluation on a commercial 5G SA network.
@@ 1. Introduction
Cellular operators could provide such monitoring solutions (cherian_lte_2013, ; margolies_can_2017, ). However, several reasons make this impractical: (1) Accuracy and Cost: Under the current LTE specification (noauthor_lte_nodate, ; noauthor_stage_nodate, ), operators cannot monitor a user’s connection with the precision required in these situations (cherian_lte_2013, ; mahyuddin_overview_2017, ), especially in NLoS scenarios involving many indoor locations. Improving precision around sensitive areas would require deploying a much denser base-station infrastructure than exists today. (2) Latency: No 3GPP-specified solution provides a geofencing result based solely on initial uplink messages; localization happens later in the connection. Without a specification change, an operator aiming to prevent communication from sensitive areas would need to block all communication until geofencing results arrive. (3) Feasibility: Every operator in the country would need to deploy this solution and collaborate with sensitive-area administrators. If even one operator refuses or is unable to cooperate, an adversary could bypass the system through that operator. Such nationwide coordination seems infeasible in practice.
On the other hand, state-of-the-art research (kotuliak_ltrack_2022, ; oh_enabling_2024, ; kumar_lte_2014, ) explores operator-independent monitoring but faces practical limitations. (1) None are shown to be low-latency when deployed in a distributed system. (2) They are restricted to a single cell per SDR, whereas realistic deployments involve multiple cells across different operators. (3) They combine downlink and uplink receivers into a single device, limiting their use to locations with good downlink visibility. These challenges make it difficult to collect high-quality signal metrics and integrate them with geofencing systems. (4) They do not support 5G SA.
@@
  • We build a new data collection platform for collecting connection data based on COTS modems and collect a large dataset of >170,000 connections with all the signal features for the initial UL messages. We make this dataset public.
@@ 2.2. Physical Uplink Channels
On the physical layer, a UE sends a message in a resource grid consisting of OFDM symbols with multiple subcarriers in each symbol. The subcarrier spacing is 15kHz for all the monitored FDD bands, and 30kHz for the 5G TDD band. For 15kHz subcarrier spacing, 14 OFDM symbols form a 1 millisecond long subframe, and 10 subframes form one frame. For 30kHz subcarrier spacing, one subframe consists of 28 OFDM symbols instead. One physical resource block (PRB) spans 12 subcarriers over 7 OFDM symbols.
@@ 3. System Design
Figure 3. In (A), we show a sketch of a deployment with two downlink receivers and six uplink receivers positioned around the sensitive area. The purple triangles on the uplink receivers show the azimuth of the two opposite-facing antennas. In (B), we show how the system components interact. First, the downlink receivers detect a new connection on the downlink and inform the UL receivers about it. The UL receivers then measure the corresponding uplink signals. Finally, all measurements are aggregated in the central unit (CU), which makes a decision about the connection. In (C), we show how a concrete wall can be used for improved inside/outside separation of a single uplink receiver unit.
@@
Accurate spatial inference about a UE connection, such as precise localization or geofencing, typically requires multiple observations of uplink messages. Given the limited number of uplink messages before connection establishment with the core network, WaveTag overcomes this constraint by deploying a distributed set of uplink receivers strategically positioned in and around the sensitive area.
@@ 3.1. System Model and Assumptions
Due to the requirement for UE detection and decision before full connection establishment with the core network, WaveTag collects metrics only for messages sent by the UE before the full connection establishment. This includes at most one PRACH preamble, one RRC Connection Request on PUSCH, and one or two subsequent PUCCH messages. Subsequent PUSCH messages after the RRC Connection Request could already include messages such as a Service Request, reestablishing an existing security context and effectively communicating with the core network.
@@ 3.3. Uplink Receivers
Uplink receivers, also operating on an SDR-based platform, are connected to two wide-beam directional antennas facing away from each other, as pictured in Figure 3. This antenna orientation enables uplink receivers to estimate the direction of the signal (see Section 5 for more details). Each SDR streams IQ samples of the uplink of one or more cellular frequency bands to an uplink receiver, where it is buffered for at least 20 milliseconds. One of the main advantages of WaveTag over existing solutions is that uplink receivers are decoupled from downlink receivers. The uplink receivers do not need to monitor downlink frequencies, and therefore they do not need to acquire a sync to the xNBs or decode downlink messages. Instead, they detect and measure uplink signals using the correct uplink allocations and timings received from the downlink receiver.
@@
All relevant downlink control messages, including uplink allocations from RARs, are sent to all uplink receivers over a low-latency communication link. Each uplink receiver then independently looks up the corresponding recorded time and frequency resources for the uplink message in the buffered IQ stream. WaveTag then extracts the signal characteristics (see Section 5) based on the reference signals and finally sends them to the central unit. Using this process, uplink receivers detect and measure PRACH, PUSCH, and PUCCH uplink messages in LTE; and PRACH and PUSCH messages in 5G.
@@ 3.4. Central Unit
Measurements of uplink messages are sent over a low-latency communication link and aggregated at the central unit. Once we receive measurements from every uplink receiver, we can further process the signal metrics for geofencing (or localization) purposes.
@@ 4. WaveTag Implementation
In this section, we explain how our system achieves both low latency and real-world scalability, including monitoring multiple cells per operator across all its allocated frequency bands. To this end, we show how the radio units acquire and process IQ samples and how the sub-components communicate with each other.
@@ 4.1. Downlink Monitoring
While downlink decoding plays a role in this process, it is not the primary contribution of this paper and is therefore not discussed in detail. Several downlink receiver implementations already exist (erni_adaptover_2022, ; erni_glados_2025, ; falkenberg_falcon_2019, ; bui_owl_2016, ; ross_towards_2023, ). We build our downlink sniffer on top of (erni_glados_2025, ) and refer the reader to that work for a detailed description of the receiver design and implementation.
@@ 4.2. Uplink Processing
Classically (kotuliak_ltrack_2022, ; hoang_ltesniffer_2023, ), uplink monitoring of an LTE cell is achieved by tuning an SDR receiver to the center frequency of the uplink of the cell and collecting 30.72 million IQ samples per second. However, in each of the frequency bands, there are three to four operators, each transmitting on different frequencies. Using the classical approach, we would require up to 24 SDRs in each uplink receiver to monitor the entire uplink spectrum. Such a system would be extremely expensive and difficult to maintain.
In our solution, we use SDRs with a sampling rate of at least 122.88 Msps, such that they have enough bandwidth to receive a whole FDD frequency band. For the 5G TDD band with cells having as much as 100 MHz bandwidth, multiple SDRs are needed. To ensure that the DC component does not disturb the uplink frequencies, we set the center frequency of our SDR between two operators allocated within a band, as close to the middle of the frequency band as possible. An uplink receiver continuously receives IQ samples from all SDRs and writes them into circular buffers. We also keep the timestamp of the last written sample. Using this timestamp, we can compute the absolute time of each IQ sample in the buffer.
@@ 4.3. Message Exchange
The downlink receiver, the uplink receivers, and the central unit continuously communicate with each other over a low-latency link. They are connected with fiber links in a star architecture with a switch in the middle. Given the relatively small number of devices, the overhead of such a network topology is negligible. In Section 6.1, we evaluate how long a message exchange takes.
@@ 5. Low-latency Geofencing
The main challenge of precise geofencing is classifying a connection’s origin—i.e., determining whether it is inside or outside a sensitive area—based on as few uplink messages as possible. To achieve this, WaveTag uses a combination of relative signal strength and signal quality features. These features are collected from multiple uplink receivers, each equipped with two directional antennas facing in opposite directions. Finally, a classification model running on the central unit infers whether the connection originates inside or outside the perimeter.
@@ 5.2. Geofencing Model
Our classification model is a fully connected neural network that uses relative features as input. We first train the model offline with Python and pyTorch (noauthor_pytorch_nodate, ). We use a simple architecture to prevent overfitting and ensure fast inference time. The architecture consists of BatchNorm, Dropout and two fully connected layers followed by a ReLU activation function. Finally, we pass the output neuron’s result through a sigmoid function to produce the classification decision. We use binary cross-entropy as a loss function and optimize the model using an Adam optimizer. We then train the model in batches and stop training once the loss no longer improves across several epochs. Finally, we perform a hyperparameter search for each deployment separately to ensure an optimal solution.
@@ 5.3. Data Collection
To improve the efficiency of the data collection, we fitted a backpack with multiple such modems that connect to the network almost every second. We can then walk a route and easily gather large amounts of uplink signal features from any location. Such an annotated dataset is crucial for training and testing of the model. We make this data available online [^1] https://doi.org/10.3929/ethz-c-000794381 to support further research.
@@ 6.1. Latency
In this section, we evaluate the latency of individual processing steps and also the end-to-end (E2E) latency of our system for various types of messages. Both individual steps and E2E latencies are shown in Figure 4. We use the Perfetto (google_googleperfetto_2024, ) and Tracy (taudul_wolfpldtracy_2025, ) libraries to measure the processing times of individual components during WaveTag’s operation. The exception is the measurement for the ZMQ message round-trip times, for which we built our own measurement tool.
@@
Downlink Decoding 540 $\mu s$ 155 $\mu s$ 47
Message Round-Trip 75 $\mu s$ 2 $\mu s$ 1000
LTE PRACH Measurement 587 $\mu s$ 147 $\mu s$ 2027
LTE PUSCH Measurement 505 $\mu s$ 129 $\mu s$ 2024
LTE PUCCH Measurement 423 $\mu s$ 145 $\mu s$ 3534
@@
LTE E2E PRACH 2223 $\mu s$ 230 $\mu s$ 48
LTE E2E PUSCH 877 $\mu s$ 128 $\mu s$ 26
LTE E2E PUCCH 1028 $\mu s$ 155 $\mu s$ 13
5G PRACH Measurement 479 $\mu s$ 51 $\mu s$ 140
5G PUSCH Measurement 602 $\mu s$ 64 $\mu s$ 140
@@
In Table 1, we show the measurement results of the individual parts of the system, as well as the E2E latencies for various message types. For the E2E PRACH measurement, we measure the elapsed time from the reception of the RAR by the downlink receiver to the completion of the inference of the geofencing model. This time encapsulates the downlink decoding of the RAR, the message exchange between the downlink receiver and the uplink receiver, the PRACH measurement by the uplink receiver, the message exchange between the uplink receiver and the central unit, and finally the model inference. For the E2E PUSCH and E2E PUCCH measurements, we measure the time from the reception of the uplink message to the inference result. This measurement does not include downlink decoding and initial message exchange, as WaveTag knows about the uplink message already 4 subframes prior to the uplink transmission.
@@ 6.3.2. Second Deployment
We trained a geofencing model as a fully connected neural network model, and furthermore, we trained a logistic regression that aggregates results over multiple messages. The estimated weights from the logistic regression are: 0.39 for PRACH, 0.38 for PUSCH and 0.23 for PUCCH, approximately matching the weights from the first deployment. The accuracy of the neural network model is 99.02% on individual messages. The ensemble model then increases the accuracy for the entire connection to 99.66%. In Table 3, we report more detailed results for the final geofencing model. Throughout the deployment, we also conducted regular spot-checks with a large number of commercial phones. We did not observe any degradation in performance for any particular model. In Appendix B, we also show how robust our model is under various input constraints.
6.4. 5G SA Feasibility Study
@@
This experiment confirms that feature extraction also works reliably in 5G SA. Even at the farthest measurement points from the uplink receiver (~100 m), we observed an SNR exceeding 20 dB. In addition, Figure 7 shows a clear spatial separation between measurement points on the left and right sides of the field. This indicates that the relative signal strength measured at the two antenna ports correlates with the direction of the uplink transmission. Given the strong similarities between the LTE and 5G PHY layers, together with the successful field evaluation, we conclude that if 5G SA were deployed at one of our sites, WaveTag would also be able to reliably geofence 5G SA connections.
@@ 7. Related Work
Other operator-independent monitoring systems that rely on uplink measurements include (kotuliak_ltrack_2022, ; oh_enabling_2024, ; kumar_lte_2014, ). They couple uplink and downlink processing on a single host using one (or more) SDRs, which leads to two drawbacks: (1) Colocation of downlink and uplink sniffers, forcing placement where good downlink reception is available and preventing flexible uplink deployment (e.g., indoors or in NLoS environments); and (2) a one-to-one mapping between uplink and downlink receivers, increasing hardware and processing requirements. In contrast, our system decouples uplink and downlink receivers, enabling independent placement and improved scalability, as one downlink receiver can serve multiple uplink receivers.
@@
The target use cases also differ. Prior systems focus on accurate localization over an active connection (e.g., looking for an uncooperative device from a moving vehicle). Because of their use cases, they can rely on a large number of PUSCH observations, rotational antennas, or even active power-boosting attacks (oh_enabling_2024, ). As a result, localization in (oh_enabling_2024, ) takes around 2 minutes, (kotuliak_ltrack_2022, ) measures ToA values of all messages from the whole connection, and (kumar_lte_2014, ) requires offline processing of recorded samples. Even though they achieve their respective tasks, these points show that they are unsuitable for low-latency geofencing. WaveTag compensates for the low number of observations with a higher number of uplink receivers, while also measuring PUCCH and PRACH messages.